# Correlator with tensor currents and two masses at two loops

### v1

@@ 4 Summary and conclusion
We have presented the first complete analytical calculation of the vacuum-to-vacuum correlator of two quark tensor currents with two massive quarks, retaining the full momentum dependence. Our results extend previous studies by providing NLO corrections for arbitrary $q^{2}$, thereby filling a gap in the literature and correcting an inconsistency in the results of Ref. [6], which seemingly omitted the renormalization of the tensor current.
@@
Our results show good agreement with a recent lattice QCD calculation [7], providing an independent cross-check and enabling a combination of perturbative and non-perturbative determinations to further improve precision. Finally, the magnitude of the remaining uncertainties motivates the calculation of NNLO corrections, which would be the natural next step toward fully controlled predictions for tensor-current correlators.
@@ Acknowledgements
We thank the Cambridge Pheno Working Group and for insightful discussions. NG is grateful to D. van Dyk for valuable input on the ideas underlying this work. NG is thanks the authors of Ref. [6] for providing their results in a machine readable form and for helpful discussion.

### v2

@@ 4 Summary and conclusion
We have presented the first complete analytical calculation of the vacuum-to-vacuum correlator of two flavour-non-diagonal tensor currents of massive quarks, retaining the full dependence on the two quark masses and the external momentum. Our results extend previous studies by providing NLO corrections for arbitrary $q^{2}$, thereby filling a gap in the literature and correcting an inconsistency in the results of Ref. [8], which seemingly omitted the renormalisation of the tensor current.
@@
Our results show good agreement with a recent lattice QCD calculation [34], providing an independent cross-check and enabling a combination of perturbative and non-perturbative determinations to further improve precision. Finally, the magnitude of the remaining uncertainties motivates the calculation of NNLO corrections, which would be the natural next step towards fully controlled predictions for tensor-current correlators.
@@ Acknowledgements
We thank the Cambridge Pheno Working Group for insightful discussions. NG is grateful to D. van Dyk for valuable input on the ideas underlying this work. NG thanks the authors of Ref. [8] for providing their results in a machine-readable form and for helpful discussion.